# AffectGPT: A New Dataset, Model, and Benchmark for Emotion Understanding with Multimodal Large Language Models

### v1

@@ Abstract
The emergence of multimodal large language models (MLLMs) advances multimodal emotion recognition (MER) to the next level—from naive discriminative tasks to complex emotion understanding with advanced video understanding abilities and natural language description. However, the current community suffers from a lack of large-scale datasets with intensive, descriptive emotion annotations, as well as a multimodal-centric framework to maximize the potential of MLLMs for emotion understanding. To address this, we establish a new benchmark for MLLM-based emotion understanding with a novel dataset (MER-Caption), and a new model (AffectGPT). Utilizing our model-based crowd-sourcing data collection strategy, we construct the largest descriptive emotion dataset to date (by far), featuring over 2K fine-grained emotion categories across 115K samples. We also introduce the AffectGPT model, designed with pre-fusion operations to enhance multimodal integration. Finally, we present MER-UniBench, a unified benchmark with evaluation metrics tailored for both typical MER tasks and the free-form, natural language output style of MLLMs. Extensive experimental results demonstrate AffectGPT’s robust performance across various MER tasks. We are publicly releasing both the AffectGPT model and the MER-Caption dataset to foster further research and development in emotion understanding.
@@ 1 Introduction
Emotions encapsulate human intentions and accurately recognizing emotional states is essential for enhancing human-computer interaction experiences [1]. Emotions can be conveyed through various human behaviors in different forms, giving rise to the task of multimodal emotion recognition (MER), which integrates multimodal information (e.g., audio, video, and text) to evaluate human emotional states. As a critical area in artificial intelligence, MER has broad applications, ranging from education [2] and psychological counseling [3] to empathic embodied robots [4].
Traditional methods primarily rely on discriminative models that map human emotions to the most likely categories from predefined emotion taxonomies. The most widely used taxonomy is Ekman’s theory [5], which classifies all emotions into six basic categories: sadness, happiness, fear, anger, surprise, and disgust. However, such categorical frameworks exhibit some limitations in modeling human affective states. For example, our emotional expressions are diverse and nuanced due to culture-specific idioms [6], context-dependent metaphors [7], and highly personalized behavioral patterns [8]. Current closed-set classification paradigms fail to capture the rich diversity of emotional expressions in real-world scenarios [9]. Meanwhile, the rigid emotion taxonomies oversimplify the continuous spectrum of emotional experiences by forcing discrete labels (e.g., anger or surprise) onto nuanced affective states that often coexist [10]. Illustrations are provided in Figure 1, where the diverse and coexisting issues are presented in Fig. (a) and (b).
Figure 1: Emotion complexity analysis. Human emotions are often diverse and coexisting. Such complex emotional states are difficult to describe using discriminative frameworks. However, MLLMs can generate emotional descriptions, offering new possibilities for complex emotion modeling. Since the original videos contain real people, to address copyright concerns, we first use DemoAI to remove personal information and then proceed with visualization.
Recent advances in multi-modal large language models (MLLMs) enable emotion understanding to move beyond traditional discriminative approaches, embracing a more generative framework [11]. This shift allows models to describe complex, coexisting emotional states in natural language. With the vast vocabulary, MLLMs can generate diverse, descriptive emotion categories beyond basic emotions, offering new opportunities for emotional understanding. However, recent research highlights that MLLMs still face limitations in emotion understanding [12, 13]. To address these challenges, this paper aims to advance emotional understanding from two key perspectives: the dataset and the model. Finally, we establish a unified benchmark tailored to the free-form, natural language output style of MLLMs.
@@
The current community still suffers from a lack of large-scale datasets with intensive, descriptive emotion annotations to realize the potential of MLLMs. The annotation strategies for constructing descriptive emotion datasets can be classified into three types: model-based, human-based, and human-model collaborative strategies. The human-based strategy is the most common way to construct emotion datasets with rich descriptive annotations. However, it’s costly to conduct crowd-sourcing to scale up the dataset size with this purely manual annotation manner. Besides, humans tend to focus on main cues, resulting in brief and incomplete descriptions [14]. Thus, researchers propose model-based automatic annotation approaches. However, due to the lack of human proofreading, this approach may result in insufficient label quality [15]. Recently, [12] propose a human-model collaborative strategy, in which models provide pre-labeled cues and humans conduct multiple rounds of checks, which can be seen as a human-led, model-assisted strategy. Although this approach offers more comprehensive descriptions, it is costly and difficult to scale the dataset. To balance label quality and dataset size, we introduce a novel annotation strategy that conducts model-based crowd-sourcing labeling with human priors, named model-led human-assisted, to construct a large-scale emotion descriptive dataset with diverse emotional categories.
@@
Categorical Dataset RAF-DB [16] I 29,672 7 Human
AffectNet [17] I 450,000 8 Human
EmoDB [18] A 535 7 Human
MSP-Podcast [19] A 73,042 8 Human
DFEW [20] V 11,697 7 Human
FERV39k [21] V 38,935 7 Human
MER2023 [22] A,V,T 5,030 6 Human
MELD [23] A,V,T 13,708 7 Human
Descriptive Dataset EmoVIT [24] I 51,200 988 Model
MERR-Coarse [15] A,V,T 28,618 113 Model
MAFW [14] A,V,T 10,045 399 Human
OV-MERD [12] A,V,T 332 236 Human-led+Model-assisted
MERR-Fine [15] A,V,T 4,487 484 Human-led+Model-assisted
@@ 2 MER-Caption: Dataset Construction
Based on the annotation manner, descriptive datasets can be categorized into model-based, human-based, and human-model collaborative strategies (see Table 1). Although the model-based approach makes it easy to expand the dataset size, it mainly relies on experience to select models and lacks human intervention, resulting in insufficient label quality [15]. To enhance label quality, [14] relied on human annotators to generate emotion descriptions. However, humans tend to focus on primary clues, easily leading to incomplete descriptions. To this end, [12] proposed a human-led, model-assisted strategy. Specifically, the model first provides pre-labeled descriptions, and then multiple annotators perform multi-round checks. Although this strategy produces more comprehensive descriptions, it comes with high annotation costs and faces challenges in scaling the dataset. In this paper, we review these annotation methods and introduce a model-led, human-assisted strategy. As shown in Figure 2, we leverage human priors to guide description generation and sample filtering, ultimately achieving automatic annotation for unlabeled data. Using this strategy, we construct the MER-Caption dataset, which includes 115K coarse-labeled samples and 31K fine-labeled samples, making a significant contribution to current descriptive datasets. Appendix H provides more comparisons with existing datasets.
@@ 2.1 Description Generation
The choice of base models is critical for generating accurate descriptions. Unlike previous work that relied solely on experience [15], we guide model selection using human priors. Specifically, we first select a small subset of samples for preliminary experiments. In this phase, we annotate fine-grained labels for each sample, allowing annotators to assign any emotions they deem appropriate, thus providing more diverse and precise labels. Based on the results in preliminary experiments (see Appendix F), we employ SALMONN [25] as the audio LLM (ALLM) to generate audio cues, Chat-UniVi [26] as the video LLM (VLLM) to extract visual cues, and GPT-3.5 [27] (“gpt-3.5-turbo-16k-0613”) to merge the audio and video cues with text content. To further reduce annotation costs, we experimented with replacing GPT-3.5 with other open-source LLMs but observed a drop in performance. The primary reason is that multimodal fusion in MER is inherently complex, often encountering issues such as modality conflict, where inconsistencies or contradictions arise between different modalities (see Appendix G). This places high demands on the LLM’s reasoning capabilities. Then, we adopt the above strategy for automatic annotation and create the MER-Caption dataset.
@@ 2.2 Sample Filtering
First, we observe that some samples contain mismatched audio and video. As shown in Figure 2, the visible person is not speaking, while the audio comes from an invisible person. This setup differs from our task, where we aim to analyze a person’s emotions based on their audio, video, and text content. Mismatched data complicates this task, shifting the focus to understanding how the interlocutor’s actions may influence the target person’s emotions. Therefore, we remove such data and plan to address this issue in future work. To automatically determine whether the visible person is speaking, we use TalkNet [28]. Preliminary experiments indicate that this tool achieves over 90% accuracy in identifying the speaking individual. Then, we remove samples with mismatched audio and video. Second, the length distribution of the generated descriptions roughly follows a Gaussian distribution (see Figure 2). Preliminary experiments reveal that descriptions at both ends of the distribution are more likely to contain errors. For instance, when ALLM and VLLM (in Section 2.1) fail to generate responses, the resulting descriptions tend to be short. As a result, we further remove descriptions located at both ends of the distribution.
@@
In addition to low-level filtering, we propose a model-based crowdsourcing technique for high-level filtering. Specifically, we train multiple multimodal emotion and sentiment classifiers using human-annotated categorical datasets. Guided by MERBench [29], we use CLIP ViT-L [30] as the visual encoder and HUBERT-L [31] as the acoustic encoder, followed by an attention-based fusion strategy to make final emotion and sentiment predictions. These pre-trained models are then used to predict labels for unlabeled data, generating multiple predictions for each sample. To mitigate potential prediction errors, we apply majority voting to determine the final label, ensuring more reliable results. We refer to this process as model-based crowdsourcing. Alternatively, emotions and sentiments can also be predicted based on the descriptions using the strategy outlined in Appendix E. If the labels extracted from the descriptions differ from those obtained through model-based crowdsourcing, we consider these descriptions to be of low quality and remove them. Through this process, we can extract knowledge from multiple human-based datasets to guide sample selection. After applying multi-level filtering, we obtain the MER-Caption+ dataset. Table 1 presents detailed comparisons between our dataset and existing ones, highlighting that our dataset is the largest multimodal emotion description dataset with diverse emotion categories.
@@ 3.2 Pre-fusion Operation
In this module, we preserve the temporal information in the vision features $\mathbf{Z_{v}}$ and audio features $\mathbf{Z_{a}}$, and utilize Q-Former [32] for multimodal fusion. Specifically, to compress the multimodal content, we first create $K$ learnable query tokens $\mathbf{Z_{q}}\in\mathbb{R}^{K\times d}$. Then, we interact $\mathbf{Z_{q}}$ with the concatenated $\mathbf{Z_{v}}$ and $\mathbf{Z_{a}}$ through cross-attention, thereby distilling the knowledge from the multimodal content into the query tokens. Formally, this process can be represented as:
@@
Unlike Q-Former which preserves temporal information, we propose a simpler architecture that directly compresses temporal information and applies attention mechanisms for multimodal fusion. This simplified module is inspired by MERBench [29], which proves that in MER tasks, features with temporal information do not always lead to better performance than compressed features. Formally, we first apply average pooling to compress unimodal features. Then, we calculate the attention weights to emphasize important modalities:
@@ A V T MER2023 MER2024 MELD IEMOCAP MOSI MOSEI SIMS SIMS v2 OV-MERD+
This task enables the prediction of fine-grained emotions, extending beyond basic categories. OV-MERD [12] is a typical dataset for this task. To improve the reliability of the evaluation results, we expand its dataset size, referring to it as OV-MERD+. For the evaluation metrics, we draw inspiration from previous work [12] and calculate results in two steps: eliminating the impact of synonyms and using set-level metrics. First, we apply a three-level grouping strategy to mitigate the impact of synonyms:
@@
  • Level 3. Emotion wheel provides natural grouping information, with core emotions displayed in the inner part and more nuanced labels in the outer part [9]. Since there is no consensus on the emotion wheel, we use $K$ emotion wheels (see Appendix K). For each sector of the emotion wheel $w_{k},k\in[1,K]$, we map all outer labels to the corresponding inner labels. This mapping function is denoted as $F_{l_{3}}^{w_{k}}(\cdot)$.
@@
Since $\mbox{F}_{\mbox{s}}$ considers both accuracy and completeness, we use it as the primary metric, with $\mbox{Precision}_{\mbox{s}}$ and $\mbox{Recall}_{\mbox{s}}$ serving as secondary metrics. To extract the predicted emotions $\mathbf{\hat{Y}}_{i}$, we employ the strategy mentioned in Appendix E.
@@
This task is a key branch of MER, whose main goal is to select the most likely label from a fixed set of basic emotions. For this task, we select four widely used benchmark datasets: MER2023 [22], MER2024 [33], IEMOCAP [34], and MELD [23]. However, the output of MLLMs, $\mathbf{\hat{Y}}_{i}=\{\hat{y}_{i}^{j}\}_{j=1}^{\hat{n}_{i}}$, contains a variable number of labels, while the dataset only provides one true label $y_{i}$. In this case, traditional metrics (such as accuracy) are not suitable for performance evaluation. To address this, we propose a new metric, hit rate, which is set to 1 when $y_{i}\in\mathbf{\hat{Y}}_{i}$ and 0 otherwise. Meanwhile, $\mathbf{\hat{Y}}_{i}$ is in free-form, while $y_{i}$ belongs to basic emotions $\mathcal{Y}$. As a result, we may encounter cases where $\hat{y}_{i}\notin\mathcal{Y}$. To this end, we also use $G_{w_{k}}(\cdot)$ and define the metric as follows:
@@
where $\mathbb{I}[\cdot]$ is an indicator function.
@@
This task is more fundamental than the two tasks mentioned above, aiming to predict the sentiment polarity. For this task, we select four benchmark datasets: CMU-MOSI [35], CMU-MOSEI [36], CH-SIMS [37], and CH-SIMS v2 [38]. For these benchmark datasets, the original labels are floating-point values, ranging from $[-1,1]$ or $[-3,3]$. We map scores of $<0$ to negative sentiment and scores of $>0$ to positive sentiment. To extract sentiment labels from the MLLM’s output, we employ the strategy outlined in Appendix E. Following previous work [35, 36], we evaluate performance using accuracy (ACC) and weighted average F-score (WAF). Due to the inherent label imbalance, we choose WAF as the primary metric and ACC as the secondary metric.

### v2

@@ Abstract
The emergence of multimodal large language models (MLLMs) advances multimodal emotion recognition (MER) to the next level—from naive discriminative tasks to complex emotion understanding with advanced video understanding abilities and natural language description. However, the current community suffers from a lack of large-scale datasets with intensive, descriptive emotion annotations, as well as a multimodal-centric framework to maximize the potential of MLLMs for emotion understanding. To address this, we establish a new benchmark for MLLM-based emotion understanding with a novel dataset (MER-Caption) and a new model (AffectGPT). Utilizing our model-based crowd-sourcing data collection strategy, we construct the largest descriptive emotion dataset to date (by far), featuring over 2K fine-grained emotion categories across 115K samples. We also introduce the AffectGPT model, designed with pre-fusion operations to enhance multimodal integration. Finally, we present MER-UniBench, a unified benchmark with evaluation metrics tailored for typical MER tasks and the free-form, natural language output style of MLLMs. Extensive experimental results show AffectGPT’s robust performance across various MER tasks. We have released both the code and the dataset to advance research and development in emotion understanding: https://github.com/zeroQiaoba/AffectGPT.
Machine Learning, ICML
@@ 1 Introduction
Emotions encapsulate human intentions, and accurately recognizing emotional states is essential for enhancing human-computer interaction experiences (Minsky, 1988). Emotions can be conveyed through various human behaviors in different forms, giving rise to the task of multimodal emotion recognition (MER), which integrates multimodal information (e.g., audio, video, and text) to evaluate human emotional states. As a critical area in artificial intelligence, MER has broad applications, ranging from education (Schutz, 2007) and psychological counseling (Liu et al., 2021) to empathic embodied robots (Spezialetti et al., 2020).
Traditional methods primarily rely on discriminative models that map human emotions to the most likely categories from predefined emotion taxonomies. The most widely used taxonomy is Ekman’s theory (Ekman & Keltner, 1970), which classifies all emotions into six basic categories: sadness, happiness, fear, anger, surprise, and disgust. However, such categorical frameworks exhibit some limitations in modeling human affective states. For example, our emotional expressions are diverse and nuanced due to culture-specific idioms (Matsumoto, 2001), context-dependent metaphors (Kövecses, 2003), and highly personalized behavioral patterns (Izard et al., 1993). Current closed-set classification paradigms fail to capture the rich diversity of emotional expressions in real-world scenarios (Plutchik, 1980). Meanwhile, the rigid emotion taxonomies oversimplify the continuous spectrum of emotional experiences by forcing discrete labels (e.g., anger or surprise) onto nuanced affective states that often coexist (Cowen & Keltner, 2017). Illustrations are provided in Figure 1, where the diverse and coexisting issues are presented in Figs. (a) and (b).
Figure 1: Emotion complexity analysis. Human emotions are often diverse and coexist simultaneously. Such complex emotional states are difficult to describe using discriminative frameworks. However, MLLMs can generate emotional descriptions, offering new possibilities for complex emotion modeling. Since the original videos contain real people, to address copyright concerns, we first use DemoAI to remove personal information and then proceed with visualization.
Recent advances in multi-modal large language models (MLLMs) enable emotion understanding to move beyond traditional discriminative approaches, embracing a more generative framework (Liang et al., 2024). This shift allows models to describe complex, coexisting emotional states in natural language. With the vast vocabulary, MLLMs can generate diverse, descriptive emotion categories beyond basic emotions, offering new opportunities for emotional understanding. However, recent research highlights that MLLMs still face limitations in emotion understanding (Lian et al., 2024a, d). To address these challenges, this paper aims to advance emotional understanding from two key perspectives: the dataset and the model. Finally, we establish a unified benchmark tailored to the free-form, natural language output style of MLLMs.
@@
The current community still suffers from a lack of large-scale datasets with intensive, descriptive emotion annotations to realize the potential of MLLMs. The annotation strategies for constructing descriptive emotion datasets can be classified into three types: model-based, human-based, and human-model collaborative strategies. The human-based strategy is the most common way to construct emotion datasets with rich descriptive annotations. However, it’s costly to conduct crowd-sourcing to scale up the dataset size with this purely manual annotation manner. Besides, humans tend to focus on main cues, resulting in brief and incomplete descriptions (Liu et al., 2022a). Thus, researchers propose model-based automatic annotation approaches. However, due to the lack of human proofreading, this approach may result in insufficient label quality (Cheng et al., 2024). Recently, Lian et al. (2024a) propose a human-model collaborative strategy, in which models provide pre-labeled cues and humans conduct multiple rounds of checks, which can be seen as a human-led, model-assisted strategy. Although this approach offers more comprehensive descriptions, it is costly and difficult to scale the dataset. To balance label quality and dataset size, we introduce a novel annotation strategy that conducts model-based crowd-sourcing labeling with human priors, named model-led human-assisted, to construct a large-scale emotion descriptive dataset with diverse emotional categories.
@@
Categorical Dataset RAF-DB (Li et al., 2017) I 29,672 7 Human
AffectNet (Mollahosseini et al., 2017) I 450,000 8 Human
EmoDB (Burkhardt et al., 2005) A 535 7 Human
MSP-Podcast (Lotfian & Busso, 2017) A 73,042 8 Human
DFEW (Jiang et al., 2020) V 11,697 7 Human
FERV39k (Wang et al., 2022) V 38,935 7 Human
MER2023 (Lian et al., 2023) A,V,T 5,030 6 Human
MELD (Poria et al., 2019) A,V,T 13,708 7 Human
Descriptive Dataset EmoVIT (Xie et al., 2024) I 51,200 988 Model
MERR-Coarse (Cheng et al., 2024) A,V,T 28,618 113 Model
MAFW (Liu et al., 2022a) A,V,T 10,045 399 Human
OV-MERD (Lian et al., 2024a) A,V,T 332 236 Human-led+Model-assisted
MERR-Fine (Cheng et al., 2024) A,V,T 4,487 484 Human-led+Model-assisted
@@ 2 MER-Caption: Dataset Construction
Based on the annotation manner, descriptive datasets can be categorized into model-based, human-based, and human-model collaborative strategies (see Table 1). Although the model-based approach makes it easy to expand the dataset size, it mainly relies on experience to select models and lacks human intervention, resulting in insufficient label quality (Cheng et al., 2024). To enhance label quality, Liu et al. (2022a) relied on human annotators to generate emotion descriptions. However, humans tend to focus on primary clues, easily leading to incomplete descriptions. To this end, Lian et al. (2024a) proposed a human-led, model-assisted strategy. Specifically, the model first provides pre-labeled descriptions, and then multiple annotators perform multi-round checks. Although this strategy produces more comprehensive descriptions, it comes with high annotation costs and faces challenges in scaling the dataset. In this paper, we review these annotation methods and introduce a model-led, human-assisted strategy. As shown in Figure 2, we leverage human priors to guide description generation and sample filtering, ultimately achieving automatic annotation for unlabeled data. Using this strategy, we construct the MER-Caption dataset, which includes 115K coarse-labeled samples and 31K fine-labeled samples, making a significant contribution to current descriptive datasets. The raw data in MER-Caption is sourced from the unlabeled portions of MER2024 (Lian et al., 2024b), with explicit permission from the dataset owners. Therefore, this paper does not involve the collection of new data but provides additional annotations for existing datasets. Appendix H provides more comparisons with existing datasets.
@@ 2.1 Description Generation
The choice of base models is critical for generating accurate descriptions. Unlike previous work that relied solely on experience (Cheng et al., 2024), we guide model selection using human priors. Specifically, we first select a small subset of samples for preliminary experiments. In this phase, we annotate fine-grained labels for each sample, allowing annotators to assign any emotions they deem appropriate, thus providing more diverse and precise labels. Based on the results in preliminary experiments (see Appendix F), we employ SALMONN (Tang et al., 2023) as the audio LLM (ALLM) to generate audio cues, Chat-UniVi (Jin et al., 2024) as the video LLM (VLLM) to extract visual cues, and GPT-3.5 (OpenAI, 2022) (“gpt-3.5-turbo-16k-0613”) to merge the audio and video cues with text content. Then, to further reduce annotation costs, we experimented with replacing GPT-3.5 with other open-source LLMs but observed a drop in performance. The primary reason is that multimodal fusion in MER is inherently complex, often encountering issues such as modality conflict, where inconsistencies or contradictions arise between different modalities (see Appendix G). This places high demands on the LLM’s reasoning capabilities. Then, we adopt the above strategy for automatic annotation and create the MER-Caption dataset.
@@ 2.2 Sample Filtering
First, we observe that some samples contain mismatched audio and video. As shown in Figure 2, the visible person is not speaking, while the audio comes from an invisible person. This setup differs from our task, where we aim to analyze a person’s emotions based on their audio, video, and text content. Mismatched data complicates this task, shifting the focus to understanding how the interlocutor’s actions may influence the target person’s emotions. Therefore, we remove such data and plan to address this issue in future work. To automatically determine whether the visible person is speaking, we use TalkNet (Tao et al., 2021). Preliminary experiments indicate that this tool achieves over 90% accuracy in identifying the speaking individual. Then, we remove samples with mismatched audio and video. Second, the length distribution of the generated descriptions roughly follows a Gaussian distribution (see Figure 2). Preliminary experiments reveal that descriptions at both ends of the distribution are more likely to contain errors. For instance, when ALLM and VLLM (in Section 2.1) fail to generate responses, the resulting descriptions tend to be short. As a result, we further remove descriptions located at both ends of the distribution.
@@
In addition to low-level filtering, we propose a model-based crowdsourcing technique for high-level filtering. Specifically, we train multiple multimodal emotion and sentiment classifiers using human-annotated categorical datasets. Guided by MERBench (Lian et al., 2024c), we use CLIP ViT-L (Radford et al., 2021) as the visual encoder and HUBERT-L (Hsu et al., 2021) as the acoustic encoder, followed by an attention-based fusion strategy to make final emotion and sentiment predictions. These pre-trained models are then used to predict labels for unlabeled data, generating multiple predictions for each sample. To mitigate potential prediction errors, we apply majority voting to determine the final label, ensuring more reliable results. We refer to this process as model-based crowdsourcing. Alternatively, emotions and sentiments can also be predicted based on the descriptions using the strategy outlined in Appendix E. If the labels extracted from the descriptions differ from those obtained through model-based crowdsourcing, we consider these descriptions to be of low quality and remove them. Through this process, we can extract knowledge from multiple human-based datasets to guide sample selection. After applying multi-level filtering, we obtain the MER-Caption+ dataset. Table 1 presents detailed comparisons between our dataset and existing ones, highlighting that our dataset is the largest multimodal emotion description dataset with diverse emotion categories.
@@ 3.2 Pre-fusion Operation
In this module, we preserve the temporal information in the vision features $\mathbf{Z_{v}}$ and audio features $\mathbf{Z_{a}}$, and utilize Q-Former (Li et al., 2023b) for multimodal fusion. Specifically, to compress the multimodal content, we first create $K$ learnable query tokens $\mathbf{Z_{q}}\in\mathbb{R}^{K\times d}$. Then, we interact $\mathbf{Z_{q}}$ with the concatenated $\mathbf{Z_{v}}$ and $\mathbf{Z_{a}}$ through cross-attention, thereby distilling the knowledge from the multimodal content into the query tokens. Formally, this process can be represented as:
@@
Unlike Q-Former which preserves temporal information, we propose a simpler architecture that directly compresses temporal information and applies attention mechanisms for multimodal fusion. This simplified module is inspired by MERBench (Lian et al., 2024c), which proves that in MER tasks, features with temporal information do not always lead to better performance than compressed features. Formally, we first apply average pooling to compress unimodal features. Then, we calculate the attention weights to emphasize important modalities:
@@
Regarding computational efficiency, the pre-fusion operation relies on Q-Former or attention mechanisms, which are significantly less computationally intensive than LLMs. Theoretically, the Q-Former enables cross-modal interaction by distilling multimodal content into query tokens, whereas the attention mechanism achieves this by dynamically computing attention weights based on multimodal inputs.
@@ A V T MER2023 MER2024 MELD IEMOCAP MOSI MOSEI SIMS SIMS v2 OV-MERD+
This task enables the prediction of fine-grained emotions, extending beyond basic categories. OV-MERD (Lian et al., 2024a) is a typical dataset for this task. To improve the reliability of the evaluation results, we expand its dataset size, referring to it as OV-MERD+. For the evaluation metrics, we draw inspiration from previous work (Lian et al., 2024a) and calculate results in two steps: eliminating the impact of synonyms and using set-level metrics. First, we apply a three-level grouping strategy to mitigate the impact of synonyms:
@@
  • Level 3. Emotion wheel provides natural grouping information, with core emotions displayed in the inner part and more nuanced labels in the outer part (Plutchik, 1980). Since there is no consensus on the emotion wheel, we use $K$ emotion wheels (see Appendix K). For each sector of the emotion wheel $w_{k},k\in[1,K]$, we map all outer labels to the corresponding inner labels. This mapping function is denoted as $F_{l_{3}}^{w_{k}}(\cdot)$.
@@
Here, $\mbox{Precision}_{\mbox{s}}$ indicates the number of correctly predicted labels, and $\mbox{Recall}_{\mbox{s}}$ indicates whether the prediction covers all ground truth. $\mbox{F}_{\mbox{s}}$ is a harmonic mean of two metrics. Since $\mbox{F}_{\mbox{s}}$ considers both accuracy and completeness, we use it as the primary metric, with $\mbox{Precision}_{\mbox{s}}$ and $\mbox{Recall}_{\mbox{s}}$ serving as secondary metrics. To extract the predicted emotions $\mathbf{\hat{Y}}_{i}$, we employ the strategy mentioned in Appendix E.
@@
This task is a key branch of MER, whose main goal is to select the most likely label from a fixed set of basic emotions. For this task, we select four widely used benchmark datasets: MER2023 (Lian et al., 2023), MER2024 (Lian et al., 2024b), IEMOCAP (Busso et al., 2008), and MELD (Poria et al., 2019). However, the output of MLLMs, $\mathbf{\hat{Y}}_{i}=\{\hat{y}_{i}^{j}\}_{j=1}^{\hat{n}_{i}}$, contains a variable number of labels, while the dataset only provides one true label $y_{i}$. In this case, traditional metrics (such as accuracy) are not suitable for performance evaluation. To address this, we propose a new metric, hit rate, which is set to 1 when $y_{i}\in\mathbf{\hat{Y}}_{i}$ and 0 otherwise. Considering that $\mathbf{\hat{Y}}_{i}$ is in free-form and $y_{i}$ belongs to basic emotions $\mathcal{Y}$, we may encounter cases where $\hat{y}_{i}\notin\mathcal{Y}$. To this end, we use the mapping function $G_{w_{k}}(\cdot)$ and define the metric as follows:
@@
where $\mathbb{I}[\cdot]$ is an indicator function. The motivation for this metric stems from the fact that basic emotion recognition tasks typically provide majority-voted labels $y_{i}$, which are generally reliable. However, emotion descriptions produce free-form outputs $\mathbf{\hat{Y}}_{i}$ that may contain multiple labels, including fine-grained ones beyond basic emotions. Therefore, we use the hit rate as the metric, ensuring that the basic label $y_{i}$ should be at least in $\mathbf{\hat{Y}}_{i}$.
During the design of this metric, we also explored the possibility of evaluating potentially incorrect labels in $\mathbf{\hat{Y}}_{i}$. However, the labels in $\mathbf{\hat{Y}}_{i}$ that differ from the basic label $y_{i}$ are not necessarily incorrect - they may represent some fine-grained emotions not covered by basic categories. Since basic emotion recognition tasks lack fine-grained reference labels, we have not yet established appropriate evaluation metrics for this purpose. This remains an important research direction for our future work.
@@
This task is more fundamental than the two tasks mentioned above, aiming to predict the sentiment polarity. For this task, we select four benchmark datasets: CMU-MOSI (Zadeh et al., 2017), CMU-MOSEI (Zadeh et al., 2018), CH-SIMS (Yu et al., 2020), and CH-SIMS v2 (Liu et al., 2022b). For these benchmark datasets, the original labels are floating-point values, ranging from $[-1,1]$ or $[-3,3]$. We map scores of $<0$ to negative sentiment and scores of $>0$ to positive sentiment. To extract sentiment labels from the MLLM’s output, we employ the strategy outlined in Appendix E. Following previous work (Zadeh et al., 2017, 2018), we evaluate performance using accuracy (ACC) and weighted average F-score (WAF). Due to the inherent label imbalance, we choose WAF as the primary metric and ACC as the secondary metric.